I need to find the line passing through (1,42) and (17,314).
Found 2 solutions by pvaka, venugopalramana:
Answer by pvaka(23)

First you have to find "m" (slope) (y2-y1)/(x2-x1)
m=314-42/17-1 =272/16 or 17
m=17
Then finding the equation of a line with the two points given.
(y-y1)=m(x-x1) or
(y-42)=17(x-1)
y-42 = 17x-17
y-42+42 = 17x-17+42
y= 17x+25

m=17 y = (0,25)
Hope this help

Answer by venugopalramana(3286)

I need to find the line passing through A(X1,Y1)=(1,42) and B(X2,Y2)= (17,314).
EQN.OF AB IS GIVEN BY
(Y-Y1)/(Y2-Y1)=(X-X1)/(X2-X1)
(Y-42)/(314-42)=(X-1)/(17-1)
16(Y-42)=272(X-1)
Y-42=17(X-1)
Y-17X-42+17=0
Y-17X-25=0